# Principal stratification with recurrent events truncated by a terminal event: A nested Bayesian nonparametric approach

### v2

@@ Acknowledgement
Research in this article was supported by the United States National Institutes of Health (NIH), National Heart, Lung, and Blood Institute (NHLBI, grant number R01-HL168202). All statements in this report, including its findings and conclusions, are solely those of the authors and do not necessarily represent the views of the NIH.
Appendix A On event rate estimand
When rate ratio estimand is of interest, an alternative definition is the equal-weighted event rate among the always-survivors, given by $\bar{\lambda}^{1}(t;r)/\bar{\lambda}^{0}(t;r)$, with $\bar{\lambda}^{z}(t;r)=\mathbb{E}\!\bigl[N_{i}^{z}(t)\big/T_{iN_{i}^{z}(t)}^{\,z}\mid\mathcal{AS}(r)\bigr]$. In this definition, each always-survivor is weighted equally regardless of their follow-up exposure time. This distinction matters when death or withdrawal truncates observation: early deaths generate exceedingly small denominators, inflating the variance. In contrast, the $\mathrm{SAER}(t;r)$ estimand is constructed as the exposure-weighted rate ratio pooling total events and total exposure time before division. Mathematically, we can write
$\frac{\mu^{z}(t;r)}{\tau^{z}(t;r)}=\mathbb{E}\!\left[\omega_{i}(t;r)\times\frac{N_{i}^{z}(t)}{T_{iN_{i}^{z}(t)}^{\,z}}\mid\mathcal{AS}(r)\right],\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \text{with}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \omega_{i}(t;r)=\frac{T_{iN_{i}^{z}(t)}^{\,z}}{\tau^{z}(t;r)}.$
Hence it is evident that $\mathrm{SAER}(t;r)$ downweights units who are observed for only a short period, and dynamically adjusts for differential survival, providing a direct public health interpretation in terms of events prevented per person-time. These properties render the exposure-weighted formulation more interpretable and operationally convenient than its equal-weighted counterpart for causal recurrent-event analyses in the presence of a competing risk of death.
Appendix B A potential pitfall of DP
B.1 DP for the recurrent event analysis with a terminal event
As recent non-causal descriptive analyses for recurrent processes with a terminal event, Paulon2020 jointly modeled the gap times of recurrent and terminal events using a Dirichlet process mixture model with log‑normal kernels for both the subject‑specific frailty parameter and the error term. Tian2024 used a DP prior for the scale parameter of Gumbel (extreme value distribution) error terms for the log-scaled survival time. These joint DP models of the log-transformed gap times and survival times, i.e., $Y^{z}_{ij}=\log(W^{z}_{ij})$ and $U^{z}_{i}=\log(D^{z}_{i})$, are typically specified as follows:
@@ B.1 DP for the recurrent event analysis with a terminal event
where $\overline{Y}_{i}=(Y_{i1},\ldots,Y_{iN_{i}})^{\top}$, $U_{i}$ is univariate, $K(\cdot)$ is an appropriate kernel function, and the mixing distribution $P$ is given a DP prior with scale parameter $\alpha>0$ and base measure $P_{0}$, denoted by $P\sim\mathrm{DP}(\alpha,P_{0})$. The second equation follows from the a.s. discrete nature of the DP, that is, the model reduces to a countable mixture with the mixing weights $w_{j}$ have a stick breaking prior with parameter $\alpha$ and $\xi_{j}\sim P_{0}$ independently of the $w_{j}$. This model can be alternatively described as the following hierarchical model: $U_{i}\mid\theta_{i}\sim F_{u}(\cdot\mid\phi_{i},\gamma_{i})$, $\overline{Y}_{i}\mid\theta_{i}\sim F_{\overline{y}}(\cdot\mid\theta_{i},\gamma_{i})$, $(\phi_{i},\theta_{i},\gamma_{i})\mid P\sim P$, $P\sim\mathrm{DP}(\alpha,P_{0\phi}\times P_{0\theta}\times P_{0\gamma}),$ where the atom $\xi_{i}$ is decomposed into $\phi_{i}$ and $\theta_{i}$, which are survival and recurrent parameters, respectively. Here the base measure $P_{0\phi}$ and $P_{0\theta}$ of the DP assumes independence between $\phi$ and $\theta$. We further assume that $P_{0\phi}$ and $P_{0\theta}$ are absolutely continuous, admitting densities $p_{0\phi}$ and $p_{0\theta}$. Then (11) can be written as $f\quantity(u,\overline{y}\mid P)=\sum_{j=1}^{\infty}w_{j}K(u\mid\overline{y},\phi_{j})K(\overline{y}\mid\theta_{j})$, where $K(u\mid\overline{y},\phi)$ and $K(\overline{y}\mid\theta)$ are the densities associated to $F_{u}(\cdot\mid\overline{y},\phi)$ and $F_{\overline{y}}(\cdot\mid\theta)$.
@@
While this is a simple, common approach adopted in literature for the joint analysis of recurrent events with a terminal event (Juhee2019; Paulon2020; Xu2021; Willem2022; Xu2022; Tian2024; Zehavi2025), a potential pitfall of these joint specifications is that the latent random partition induced by the DP will be overwhelmingly determined by the recurrent events rather than the terminal event as the number of recurrences grows. This issue occurs because the likelihood contribution of the recurrent event dominates that of a single terminal event, which may not even be observed due to other censoring events. Consequently, the DP prior favors a random partition that approximates the distribution of the recurrent events well with as many clusters as necessary. However, having too many small clusters, which are dominantly determined by the contribution of $\overline{y}$, can lead to unreliable within-cluster predictions of the survival event $u$, with higher posterior variance due to unnecessarily small sample sizes, thereby compromising the predictive performance for the survival event. Predictive performance is crucial for Bayesian causal inference because the inferential procedure imputes missing potential outcomes and, therefore, heavily depends on the accuracy of potential outcome predictions. Additionally, existing approaches use the joint DP primarily to model a frailty term. However, because the frailty is widely recognized as a subject-specific distinct parameter that reflects an individual’s vulnerability (e.g., physical health conditions), it is somewhat counterintuitive that the random partition, which yields the subject-specific frailty, is driven largely by numerous recurrent events rather than the subject-specific survival event.
B.2 Random partition induced by the standard DP
@@
The prior on the random partition induced by the DP is $p(\mathcal{P}_{n})\propto\alpha^{M_{n}}\prod_{j=1}^{M_{n}}\Gamma(n_{j}),$ where $n_{j}=|\mathcal{S}_{j}|$ (Antoniak1974) and $\alpha$ is a precision parameter of the DP. Consequently, the posterior distribution of the random partition $\mathcal{P}_{n}$ can be written as
@@
Appendix C Proofs
C.1 Proof of Proposition 1
@@
The result of Antoniak1974 implies that the random partition induced by the DP is $p(\mathcal{P}_{n,u})=\frac{\Gamma(\alpha_{\phi})}{\Gamma(\alpha_{\phi}+n)}\alpha_{\phi}^{M_{n}}\prod_{j=1}^{M_{n}}\Gamma(n_{j})$ and $p(\mathcal{P}_{n,y}\mid\phi_{j}^{*})=\frac{\Gamma(\alpha_{\theta}(\phi_{j}^{*}))}{\Gamma(\alpha_{\theta}(\phi_{j}^{*})+n_{j})}\alpha_{\theta}(\phi_{j}^{*})^{M_{n,j}}\prod_{l=1}^{M_{n,j}}\Gamma(n_{l\mid j})$. Putting these into (15) and integrating out $\phi^{*}$, we obtain the desired result for the first part of the proposition. The second part of the proposition can be obtained by the Bayes theorem. ∎
C.2 Proof of Theorem 1
@@
$\mathrm{Pr}\left(N_{i}^{z}(t)=n\mid\mathcal{AS}(r)\right)$ $=\mathrm{Pr}\left(N_{i}^{z}(t)=n\mid T^{1}_{iN_{i}^{1}(r)}<D_{i}^{1},T_{iN_{i}^{0}(r)}^{0}<D_{i}^{0}\right)$
$=\mathbb{E}_{X,\gamma\mid\mathcal{AS}(r)}\left[\mathrm{Pr}\left(N_{i}^{z}(t)=n\mid T^{1}_{iN_{i}^{1}(r)}<D_{i}^{1},T_{iN_{i}^{0}(r)}^{0}<D_{i}^{0},X,\gamma\right)\right]$
$=E_{X,\gamma\mid\mathcal{AS}(r)}\quantity[\frac{\mathrm{Pr}\quantity(N_{i}^{z}(t)=n,T^{1}_{iN_{i}^{1}(r)}<D_{i}^{1},T_{iN_{i}^{0}(r)}^{0}<D_{i}^{0}\mid X,\gamma)}{\mathrm{Pr}\quantity(T^{1}_{iN_{i}^{1}(r)}<D_{i}^{1},T_{iN_{i}^{0}(r)}^{0}<D_{i}^{0}\mid X,\gamma)}]$
$=E_{X,\gamma\mid\mathcal{AS}(r)}\quantity[\frac{\mathrm{Pr}\quantity(N_{i}^{z}(t)=n,T^{z}_{iN_{i}^{z}(r)}<D_{i}^{z}\mid X,\gamma)\mathrm{Pr}\quantity(T_{iN_{i}^{1-z}(r)}^{z}<D_{i}^{1-z}\mid X,\gamma)}{\mathrm{Pr}\quantity(T^{1}_{iN_{i}^{1}(r)}<D_{i}^{1}\mid X,\gamma)\mathrm{Pr}\quantity(T_{iN_{i}^{0}(r)}^{0}<D_{i}^{0}\mid X,\gamma)}]$
$=E_{X,\gamma\mid\mathcal{AS}(r)}\quantity[\mathrm{Pr}\quantity(N_{i}^{z}(t)=n\mid T^{z}_{iN_{i}^{z}(r)}<D_{i}^{z},X,\gamma)]$
$=E_{X,\gamma\mid\mathcal{AS}(r)}\quantity[\mathrm{Pr}\quantity(N_{i}^{z}(t)=n\mid Z_{i}=z,T^{z}_{iN_{i}^{z}(r)}<D_{i}^{z},X,\gamma)]$
$=E_{X,\gamma\mid\mathcal{AS}(r)}\quantity[\mathrm{Pr}\quantity(N_{i}(t)=n\mid Z_{i}=z,T_{iN_{i}(r)}<D_{i},X,\gamma)]$
$=\int_{\mathcal{X}}\int_{0}^{\infty}\mathrm{Pr}\quantity(N_{i}(t)=n\mid Z_{i}=z,T_{iN_{i}(r)}<D_{i},X,\gamma)f_{X,\gamma\mid\mathcal{AS}(r)}(x,\gamma)d\gamma dx.$
@@
$f_{X,\gamma\mid\mathcal{AS}(r)}(x,\gamma)=\frac{\mathrm{Pr}\quantity(T^{1}_{iN_{i}^{1}(r)}<D_{i}^{1},T_{iN_{i}^{0}(r)}^{0}<D_{i}^{0}\mid\mathbf{X}_{i}=\mathbf{x},\gamma)f_{\gamma}(\gamma)f_{X}(x)}{\int_{\mathcal{X}}\int_{0}^{\infty}\mathrm{Pr}\quantity(T^{1}_{iN_{i}^{1}(r)}<D_{i}^{1},T_{iN_{i}^{0}(r)}^{0}<D_{i}^{0}\mid\mathbf{X}_{i}=\mathbf{x}^{\prime},\gamma^{\prime})f_{\gamma}(\gamma^{\prime})f_{X}(x^{\prime})d\gamma^{\prime}dx^{\prime}}$
$=\frac{\mathrm{Pr}\quantity(T^{1}_{iN_{i}^{1}(r)}<D_{i}^{1}\mid\mathbf{X}_{i}=\mathbf{x},\gamma)\mathrm{Pr}\quantity(T_{iN_{i}^{0}(r)}^{0}<D_{i}^{0}\mid\mathbf{X}_{i}=\mathbf{x},\gamma)f_{\gamma}(\gamma)f_{X}(x)}{\int_{\mathcal{X}}\int_{0}^{\infty}\mathrm{Pr}\quantity(T^{1}_{iN_{i}^{1}(r)}<D_{i}^{1}\mid\mathbf{X}_{i}=\mathbf{x}^{\prime},\gamma^{\prime})\mathrm{Pr}\quantity(T_{iN_{i}^{0}(r)}^{0}<D_{i}^{0}\mid\mathbf{X}_{i}=\mathbf{x}^{\prime},\gamma^{\prime})f_{\gamma}(\gamma^{\prime})f_{X}(x^{\prime})d\gamma^{\prime}dx^{\prime}}$
$=\frac{\mathrm{Pr}\quantity(T^{1}_{iN_{i}^{1}(r)}<D_{i}^{1}\mid Z_{i}=1,\mathbf{X}_{i}=\mathbf{x},\gamma)\mathrm{Pr}\quantity(T_{iN_{i}^{0}(r)}^{0}<D_{i}^{0}\mid Z_{i}=0,\mathbf{X}_{i}=\mathbf{x},\gamma)f_{\gamma}(\gamma)f_{X}(x)}{\int_{\mathcal{X}}\int_{0}^{\infty}\mathrm{Pr}\quantity(T^{1}_{iN_{i}^{1}(r)}<D_{i}^{1}\mid Z_{i}=1,\mathbf{X}_{i}=\mathbf{x}^{\prime},\gamma^{\prime})\mathrm{Pr}\quantity(T_{iN_{i}^{0}(r)}^{0}<D_{i}^{0}\mid Z_{i}=0,\mathbf{X}_{i}=\mathbf{x}^{\prime},\gamma^{\prime})f_{\gamma}(\gamma^{\prime})f_{X}(x^{\prime})d\gamma^{\prime}dx^{\prime}}$
$=\frac{\mathrm{Pr}\quantity(T_{iN_{i}(r)}<D_{i}\mid Z_{i}=1,\mathbf{X}_{i}=\mathbf{x},\gamma)\mathrm{Pr}\quantity(T_{iN_{i}(r)}<D_{i}\mid Z_{i}=0,\mathbf{X}_{i}=\mathbf{x},\gamma)f_{\gamma}(\gamma)f_{X}(x)}{\int_{\mathcal{X}}\int_{0}^{\infty}\mathrm{Pr}\quantity(T_{iN_{i}(r)}<D_{i}\mid Z_{i}=1,\mathbf{X}_{i}=\mathbf{x}^{\prime},\gamma^{\prime})\mathrm{Pr}\quantity(T_{iN_{i}(r)}<D_{i}\mid Z_{i}=0,\mathbf{X}_{i}=\mathbf{x}^{\prime},\gamma^{\prime})f_{\gamma}(\gamma^{\prime})f_{X}(x^{\prime})d\gamma^{\prime}dx^{\prime}}$
$=\frac{\eta_{r}(1,x,\gamma)\eta_{r}(0,x,\gamma)f_{\gamma}(\gamma)f_{X}(x)}{\int_{\mathcal{X}}\int_{0}^{\infty}\eta_{r}(1,x,\gamma)\eta_{r}(0,x,\gamma)f_{\gamma}(\gamma^{\prime})f_{X}(x^{\prime})d\gamma^{\prime}dx^{\prime}},$
where the first line follows from Bayes’ theorem, the second line follows from Assumption 4, the third line follows from Assumption 2, and the fourth line follows from Assumption 1. It is easy to show that (2) can be identified in the same manner. ∎
Appendix D Details of Gibbs sampler

### v3

@@ Acknowledgement
Research in this article was supported by the United States National Institutes of Health (NIH), National Heart, Lung, and Blood Institute (NHLBI, grant number R01-HL168202 and 1R01HL178513).
Appendix A A potential pitfall of DP
A.1 DP for the recurrent event analysis with a terminal event
As recent non-causal descriptive analyses for recurrent processes with a terminal event, Paulon et al. (2020) jointly modeled the gap times of recurrent and terminal events using a Dirichlet process mixture model with log-normal kernels for both the subject-specific frailty parameter and the error term. Tian et al. (2024) used a DP prior for the scale parameter of Gumbel (extreme value distribution) error terms for the log-scaled survival time. These joint DP models of the log-transformed gap times and survival times, i.e., $Y^{z}_{ij}=\log(W^{z}_{ij})$ and $U^{z}_{i}=\log(D^{z}_{i})$, are typically specified as follows:
@@ A.1 DP for the recurrent event analysis with a terminal event
where $\overline{Y}_{i}=(Y_{i1},\ldots,Y_{iN_{i}})^{\top}$, $U_{i}$ is univariate, $K(\cdot)$ is an appropriate kernel function, and the mixing distribution $P$ is given a DP prior with scale parameter $\alpha>0$ and base measure $P_{0}$, denoted by $P\sim\mathrm{DP}(\alpha,P_{0})$. The second equation follows from the a.s. discrete nature of the DP, that is, the model reduces to a countable mixture with the mixing weights $w_{j}$ having a stick breaking prior with parameter $\alpha$ and $\xi_{j}\sim P_{0}$ independently of the $w_{j}$. This model can be alternatively described as the following hierarchical model: $U_{i}\mid\theta_{i}\sim F_{u}(\cdot\mid\phi_{i},\gamma_{i})$, $\overline{Y}_{i}\mid\theta_{i}\sim F_{\overline{y}}(\cdot\mid\theta_{i},\gamma_{i})$, $(\phi_{i},\theta_{i},\gamma_{i})\mid P\sim P$, $P\sim\mathrm{DP}(\alpha,P_{0\phi}\times P_{0\theta}\times P_{0\gamma}),$ where the atom $\xi_{i}$ is decomposed into $\phi_{i}$ and $\theta_{i}$, which are survival and recurrent parameters, respectively. Here the base measure $P_{0\phi}$ and $P_{0\theta}$ of the DP assumes independence between $\phi$ and $\theta$. We further assume that $P_{0\phi}$ and $P_{0\theta}$ are absolutely continuous, admitting densities $p_{0\phi}$ and $p_{0\theta}$. Then (9) can be written as $f\quantity(u,\overline{y}\mid P)=\sum_{j=1}^{\infty}w_{j}K(u\mid\overline{y},\phi_{j})K(\overline{y}\mid\theta_{j})$, where $K(u\mid\overline{y},\phi)$ and $K(\overline{y}\mid\theta)$ are the densities associated to $F_{u}(\cdot\mid\overline{y},\phi)$ and $F_{\overline{y}}(\cdot\mid\theta)$.
@@
While this is a simple, common approach adopted in literature for the joint analysis of recurrent events with a terminal event (Lee et al., 2019; Paulon et al., 2020; Xu et al., 2021; van den Boom et al., 2022; Xu et al., 2022; Tian et al., 2024; Zehavi et al., 2025), a potential pitfall of these joint specifications is that the latent random partition induced by the DP will be overwhelmingly determined by the recurrent events rather than the terminal event as the number of recurrences grows. This issue occurs because the likelihood contribution of the recurrent event dominates that of a single terminal event, which may not even be observed due to other censoring events. Consequently, the DP prior favors a random partition that approximates the distribution of the recurrent events well with as many clusters as necessary. However, having too many small clusters, which are dominantly determined by the contribution of $\overline{y}$, can lead to unreliable within-cluster predictions of the survival event $u$, with higher posterior variance due to unnecessarily small sample sizes, thereby compromising the predictive performance for the survival event. Predictive performance is crucial for Bayesian causal inference because the inferential procedure imputes missing potential outcomes and, therefore, heavily depends on the accuracy of potential outcome predictions. Additionally, existing approaches use the joint DP primarily to model a frailty term. However, because the frailty is widely recognized as a subject-specific distinct parameter that reflects an individual’s vulnerability (e.g., physical health conditions), it is somewhat counterintuitive that the random partition, which yields the subject-specific frailty, is driven largely by numerous recurrent events rather than the subject-specific survival event.
A.2 Random partition induced by the standard DP
@@
The prior on the random partition induced by the DP is $p(\mathcal{P}_{n})\propto\alpha^{M_{n}}\prod_{j=1}^{M_{n}}\Gamma(n_{j}),$ where $n_{j}=|\mathcal{S}_{j}|$ (Antoniak, 1974) and $\alpha$ is a precision parameter of the DP. Consequently, the posterior distribution of the random partition $\mathcal{P}_{n}$ can be written as
@@
Appendix B Proofs
B.1 Proof of Proposition 1
@@
The result of Antoniak (1974) implies that the random partition induced by the DP is $p(\mathcal{P}_{n,u})=\frac{\Gamma(\alpha_{\phi})}{\Gamma(\alpha_{\phi}+n)}\alpha_{\phi}^{M_{n}}\prod_{j=1}^{M_{n}}\Gamma(n_{j})$ and $p(\mathcal{P}_{n,y}\mid\phi_{j}^{*})=\frac{\Gamma(\alpha_{\theta}(\phi_{j}^{*}))}{\Gamma(\alpha_{\theta}(\phi_{j}^{*})+n_{j})}\alpha_{\theta}(\phi_{j}^{*})^{M_{n,j}}\prod_{l=1}^{M_{n,j}}\Gamma(n_{l\mid j})$. Putting these into (13) and integrating out $\phi^{*}$, we obtain that the prior on the random partition induced by the EDP is
The claim in the proposition can be obtained by the Bayes theorem. ∎
B.2 Proof of Theorem 1
@@
$\mathrm{Pr}\left(N_{i}^{z}(t)=n\mid\mathcal{AS}(r)\right)$ $=\mathrm{Pr}\left(N_{i}^{z}(t)=n\mid r<D_{i}^{1},r<D_{i}^{0}\right)$
$=\mathbb{E}_{X,\boldsymbol{\gamma}\mid\mathcal{AS}(r)}\left[\mathrm{Pr}\left(N_{i}^{z}(t)=n\mid r<D_{i}^{1},r<D_{i}^{0},X,\boldsymbol{\gamma}\right)\right]$
$=\mathbb{E}_{X,\boldsymbol{\gamma}\mid\mathcal{AS}(r)}\quantity[\frac{\mathrm{Pr}\quantity(N_{i}^{z}(t)=n,r<D_{i}^{1},r<D_{i}^{0}\mid X,\boldsymbol{\gamma})}{\mathrm{Pr}\quantity(r<D_{i}^{1},r<D_{i}^{0}\mid X,\boldsymbol{\gamma})}]$
$=\mathbb{E}_{X,\boldsymbol{\gamma}\mid\mathcal{AS}(r)}\quantity[\frac{\mathrm{Pr}\quantity(N_{i}^{z}(t)=n,r<D_{i}^{z}\mid X,\gamma^{z})\mathrm{Pr}\quantity(r<D_{i}^{1-z}\mid X,\gamma^{1-z})}{\mathrm{Pr}\quantity(r<D_{i}^{1}\mid X,\gamma^{1})\mathrm{Pr}\quantity(r<D_{i}^{0}\mid X,\gamma^{0})}]$
$=\mathbb{E}_{X,\boldsymbol{\gamma}\mid\mathcal{AS}(r)}\quantity[\mathrm{Pr}\quantity(N_{i}^{z}(t)=n\mid r<D_{i}^{z},X,\gamma^{z})]$
$=\mathbb{E}_{X,\boldsymbol{\gamma}\mid\mathcal{AS}(r)}\quantity[\mathrm{Pr}\quantity(N_{i}^{z}(t)=n\mid Z_{i}=z,r<D_{i}^{z},X,\gamma^{z})]$
$=\mathbb{E}_{X,\boldsymbol{\gamma}\mid\mathcal{AS}(r)}\quantity[\mathrm{Pr}\quantity(N_{i}(t)=n\mid Z_{i}=z,r<D_{i},X,\gamma^{z})]$
$=\int_{\mathcal{X}}\int_{\Gamma}\mathrm{Pr}\quantity(N_{i}(t)=n\mid Z_{i}=z,r<D_{i},X,\gamma^{z})f_{X,\boldsymbol{\gamma}\mid\mathcal{AS}(r)}(x,\boldsymbol{\gamma})d\boldsymbol{\gamma}dx.$
@@
$f_{X,\boldsymbol{\gamma}\mid\mathcal{AS}(r)}(x,\boldsymbol{\gamma})=\frac{\mathrm{Pr}\quantity(r<D_{i}^{1},r<D_{i}^{0}\mid\mathbf{X}_{i}=\mathbf{x},\boldsymbol{\gamma})f_{\boldsymbol{\gamma}}(\boldsymbol{\gamma})f_{X}(x)}{\int_{\mathcal{X}}\int_{\Gamma}\mathrm{Pr}\quantity(r<D_{i}^{1},r<D_{i}^{0}\mid\mathbf{X}_{i}=\mathbf{x}^{\prime},\boldsymbol{\gamma}^{\prime})f_{\boldsymbol{\gamma}}(\boldsymbol{\gamma}^{\prime})f_{X}(x^{\prime})d\boldsymbol{\gamma}^{\prime}dx^{\prime}}$
$=\frac{\mathrm{Pr}\quantity(r<D_{i}^{1}\mid\mathbf{X}_{i}=\mathbf{x},\gamma^{1})\mathrm{Pr}\quantity(r<D_{i}^{0}\mid\mathbf{X}_{i}=\mathbf{x},\gamma^{0})f_{\boldsymbol{\gamma}}(\boldsymbol{\gamma})f_{X}(x)}{\int_{\mathcal{X}}\int_{\Gamma}\mathrm{Pr}\quantity(r<D_{i}^{1}\mid\mathbf{X}_{i}=\mathbf{x}^{\prime},\gamma^{1^{\prime}})\mathrm{Pr}\quantity(r<D_{i}^{0}\mid\mathbf{X}_{i}=\mathbf{x}^{\prime},\gamma^{0^{\prime}})f_{\boldsymbol{\gamma}}(\boldsymbol{\gamma}^{\prime})f_{X}(x^{\prime})d\boldsymbol{\gamma}^{\prime}dx^{\prime}}$
$=\frac{\mathrm{Pr}\quantity(r<D_{i}^{1}\mid Z_{i}=1,\mathbf{X}_{i}=\mathbf{x},\gamma^{1})\mathrm{Pr}\quantity(r<D_{i}^{0}\mid Z_{i}=0,\mathbf{X}_{i}=\mathbf{x},\gamma^{0})f_{\boldsymbol{\gamma}}(\boldsymbol{\gamma})f_{X}(x)}{\int_{\mathcal{X}}\int_{\Gamma}\mathrm{Pr}\quantity(r<D_{i}^{1}\mid Z_{i}=1,\mathbf{X}_{i}=\mathbf{x}^{\prime},\gamma^{1^{\prime}})\mathrm{Pr}\quantity(r<D_{i}^{0}\mid Z_{i}=0,\mathbf{X}_{i}=\mathbf{x}^{\prime},\gamma^{0^{\prime}})f_{\boldsymbol{\gamma}}(\boldsymbol{\gamma}^{\prime})f_{X}(x^{\prime})d\boldsymbol{\gamma}^{\prime}dx^{\prime}}$
$=\frac{\mathrm{Pr}\quantity(r<D_{i}\mid Z_{i}=1,\mathbf{X}_{i}=\mathbf{x},\gamma^{1})\mathrm{Pr}\quantity(r<D_{i}\mid Z_{i}=0,\mathbf{X}_{i}=\mathbf{x},\gamma^{0})f_{\boldsymbol{\gamma}}(\boldsymbol{\gamma})f_{X}(x)}{\int_{\mathcal{X}}\int_{\Gamma}\mathrm{Pr}\quantity(r<D_{i}\mid Z_{i}=1,\mathbf{X}_{i}=\mathbf{x}^{\prime},\gamma^{1^{\prime}})\mathrm{Pr}\quantity(r<D_{i}\mid Z_{i}=0,\mathbf{X}_{i}=\mathbf{x}^{\prime},\gamma^{0^{\prime}})f_{\boldsymbol{\gamma}}(\boldsymbol{\gamma}^{\prime})f_{X}(x^{\prime})d\boldsymbol{\gamma}^{\prime}dx^{\prime}}$
$=\frac{\eta_{r}(1,x,\gamma^{1})\eta_{r}(0,x,\gamma^{0})f_{\boldsymbol{\gamma}}(\boldsymbol{\gamma})f_{X}(x)}{\int_{\mathcal{X}}\int_{\Gamma}\eta_{r}(1,x^{\prime},\gamma^{1^{\prime}})\eta_{r}(0,x^{\prime},\gamma^{0^{\prime}})f_{\boldsymbol{\gamma}}(\boldsymbol{\gamma}^{\prime})f_{X}(x^{\prime})d\boldsymbol{\gamma}^{\prime}dx^{\prime}},$
where the first line follows from Bayes’ theorem, the second line follows from Assumption 4, the third line follows from Assumption 2, and the fourth line follows from Assumption 1. ∎
Appendix C Details of Gibbs sampler